OpenStudy (anonymous):

Create your own binomial expression with a radical in the second term. Part 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why?

OpenStudy (anonymous):

The pupose of a conjugate is to eliminate radicals, imaginary numbers, etc., so that the expression is real-valued. So, for example, take a binomial expression like \[2-\sqrt{5}\]. The conjugate would then be \[2+\sqrt{5}.\] Multiplying the two conjugates together gives the following expession: \[(2-\sqrt{5})(2+\sqrt{5})=4+2\sqrt{5}-2\sqrt{5}-(\sqrt{5)^{2}}=4-5=-1\] I hope that helps.